OpenStudy (anonymous):

I would just like assistance with the problem, if possible. If a stone is tossed from the top of a 190 meter building, the height of the stone as a function of time is given by h(t)=-9.8t^2-10t +190, where t is in seconds, and height is in meters. After how many seconds will the stone hit the ground? Round to the nearest hundredth's place; include units in your answer. This is the help I had from another person on here. h(t)=-9.8t^2 -10t +190 190=-9.8t^2-10t+190 -9.8t^2-10t=0 t= 1.02 seconds or t=0 secs How do I check this, or what is next?

OpenStudy (bahrom7893):

okay ill check this over..

OpenStudy (bahrom7893):

-9.8t^2-10t +190=0 he was wrong.. actually..

OpenStudy (bahrom7893):

or she.. whoever helped u.. that's the right equation..

OpenStudy (bahrom7893):

-9.8t^2-10t +190=0, t=3.9, disregard the negative answer since that doesn't make sense..

OpenStudy (anonymous):

So this goes after the third line?

OpenStudy (bahrom7893):

No this goes initially.. just cross everything u did before out and write this: -9.8t^2-10t +190=0 then solve for ts and one t will be -number, and 2nd one is t=3.9

OpenStudy (anonymous):

so what do I put this into to check it?

OpenStudy (bahrom7893):

i used a calculator it is correct.

OpenStudy (anonymous):

I don't doubt it, I need to show it in an assignment. How do you make it into this (x+?)(x-?)

OpenStudy (bahrom7893):

You have to use a quadratic equation, you know D = b^2-4ac

OpenStudy (bahrom7893):

etc..

OpenStudy (anonymous):

Nickie, do you understand why -9.8t^2-10t +190 equals 0 and not 190?

OpenStudy (anonymous):

No, I do not understand why there was a zero just added to the end of the equation, usually it is you minus from one side and put it in as the opposite on the other side of the equal sign.

OpenStudy (anonymous):

okay. h(t) = -9.8t^2-10t +190 represents the height of the stone from the ground at time t correct?

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

okay, then what is the height of the stone at the instant it hits the ground?

OpenStudy (anonymous):

d=7548, is what I have.

OpenStudy (anonymous):

So for the whole problem I have: h(t)=-9.8t^2 -10t +190 -9.8t^2-10t+190=0 t=-3.9 or t=3.9 D=b^2-4ac a=-9.8 b=-10 c= 190 -10^2-4*-9.8*190 100-4*-9.8 100+39.2*190 100+7448 =7548 Is this right? So what is the final answer t=-3.9 or t=3.9? Or is it 7548?